Please help me find (a)2v + u, (b)u -3v, and (c)5u - v in the equation of u = 2i + 3j, v = -i - 2j.

Answer by ikleyn(52949)

.
u = 2i + 3j, v = -i - 2j.

Then

(a)  2v + u = (-2i - 4j) + (2i + 3j) = (-2 + 2)i + (-4 + 3)j = 0*i + (-1)*j = -j .
               --------     -------
                  2v           u 


(b)  u -3v = (2i + 3j) - (-3i - 6j) = (2 + 3)i + (3 + 6)j = 5i + 9j .
              -------     ---------
                 u            3v


(c)  5u - v = (10i + 15j) - (-i - 2j) = (10+1)i + (15 + 2)j = 11i + 17j .
               ---------     -------
                  5u            v